# The Interaction Layer: An Exploration for Co-Designing User-LLM Interactions in Parental Wellbeing Support Systems

### v1

@@ 3.1. Origins and motivations for the initial prototype
Our parental wellbeing support prototype, was based on a community wellbeing project called Nurture, initially run by a nonprofit organisation dedicated to children’s wellbeing [Org name removed for review]. The original Nurture project was a peer-to-peer mentoring initiative designed to support new mothers during their baby’s first year. In this program, experienced volunteer mothers were trained to mentor new moms, providing guidance and encouragement. A digital platform supporting this program offered personalised, week-by-week content delivered via text messages, acting as a conversation catalyst between mentors and mentees. The prompts were designed to initiate meaningful dialogue and help new moms navigate early motherhood.
@@ 3.2. Designing the initial prototype
Recognising the need for a scalable solution that could retain the core support offered by Nurture, we partnered with the organisation to collate resources and findings from Nurture project for developing a new AI tool. In collaboration with the organisation’s project team, we wrote the initial prompt for a basic large language model (LLM)-based chatbot using Open AI’s GPT-4 API (OpenAI, 2023), zero-shot prompting, which we named NurtureBot v1.
NurtureBot v1 was designed to replicate the supportive environment of the original peer-to-peer texting interactions while introducing the scalability and accessibility of AI. The chatbot was refined to deliver three key features: empathetic chatting, wellbeing exercises, and parenting information, all adapted for AI-driven interactions. Additionally, it incorporated ”conversation starters” thoughtfully crafted by Nurture mentors. Empathetic chatting was intended to provide an active listening that reflect warmth and understanding. Wellbeing exercises were included to offer practical tools for managing stress and mental wellbeing, similar to those shared by mentors. Lastly, parenting information was integrated to ensure that users could access reliable and evidence-based resources tailored to their needs, much like the information shared in the original peer support sessions.
Leveraging insights from the Nurture project in Part I, we designed NurtureBot v1 and initiated iterative trials in the following Part II by testing the prototype with everyday users.
4. Part II: NurtureBot v1 Testing and Analysis
@@
4.1. Participants and Method
Once we had secured ethical approval via our institutional ethics board [removed for review], we recruited 32 parents of children aged 5 or 6 through the Prolific platform (Prolific, 2014). The average age of participants was 41.17 years, with 20 females and 12 males. Five participants (15%) were in the habit of using ChatGPT everyday. Participants self-reported no acute stress or mental or physical health issues during the study period, nor to have they received therapy in the 12 months prior. The study’s purpose and methodology were explained to parents through an information sheet, after which informed consent was obtained. They were paid  \$3 for each day of participation.
Over the five-day period, participants engaged with NurtureBot v1 for at least five minutes each day, with each day dedicated to simulating a specific stage of early parenthood: pregnancy, the week of childbirth, few weeks postpartum, one month, and six months after birth. The objective was to assess how well the NurtureBot v1 could support them in these hypothetical scenarios. Each day, participants interacted with the chatbot, receiving responses that included links to resources, empathetic messages, and suggestions for wellbeing exercises (see Figure 3). At the end of each day, after interacting with NurtureBot hosted on a website via Streamlit (Streamlit, 2019), participants were asked to complete a survey via Qualtrics (Qualtrics, 2005), providing feedback on their experience with the chatbot by answering the questions listed in Figure 2. On the final day, they also completed the Chatbot Usability Questionnaire (CUQ), a specialised questionnaire designed to access the usability of healthcare chatbots (Holmes et al., 2019).
@@ 4.1. Participants and Method
4.2. Data Collection and Analysis
Over the course of the five-day study, 144 conversations and associated feedback were generated, denoting a daily-average of 90% (28.8 out of 32 participants). All data was securely stored on departmental servers, compliant with GDPR (Union, 2016), and no personal data was collected, consistent with Prolific platform’s design.
The qualitative analysis of the feedback from the survey (See questions 1 and 2 in Figure 2) was conducted using a thematic analysis approach (Braun and Clarke, 2012). Our analysis focused on the instances that captured parent reactions to interacting with NurtureBot. The first author conducted the initial analysis using inductive, open coding, and regularly met with the second author to discuss emerging themes and explore alternative interpretations of participant responses. While participants did share positive feedback on several aspects of the chatbot, highlighting its usefulness and relevance, especially during early stages of parenthood, we do not explore the parents perspectives on the potential of LLM-based parental wellbeing support as this is reported in the literature (Ashraf, 2024), and to keep our focus on answering RQ1. Hence, we focused on analysing 131 of the 288 (45.48%) pieces of feedback that highlighted negative interactions and suggestions for improvement. This led us to the identification of 28 unique pain points, which were further grouped into seven overarching themes, converged in three subsequent meetings. These themes highlighted the key problems participants faced while interacting with NurtureBot v1 as exemplified below. We provide only a high-level summary here to conserve the reader’s focus for the ARC section, which forms the core of the HCI discussion.

### v2

@@ 3.1. Origins and motivations for the initial prototype
Our parental wellbeing support prototype was built upon Nurture, a community wellbeing project originally run by the nonprofit organisation Committee for Children (CFC) (Committee for Children, [n. d.]), which is dedicated to promoting children’s wellbeing. The original Nurture project was a peer-to-peer mentoring initiative designed to support new mothers during their baby’s first year. In this program, experienced volunteer mothers were trained to mentor new moms, providing guidance and encouragement. A digital platform supporting this program offered personalised, week-by-week content delivered via text messages, acting as a conversation catalyst between mentors and mentees. These messages were designed to initiate meaningful dialogue and help new moms navigate early motherhood.
@@ 3.2. Designing the initial prototype
Recognising the need for a scalable solution that could retain the core support offered by Nurture, we partnered with CFC to collate resources and findings from Nurture project for developing a new AI tool. Based on our learning from the Nurture project, we developed an initial prompt for a basic LLM-based chatbot by zero-shot prompting Open AI’s GPT-4 API (OpenAI, 2023), which we named NurtureBot v1. We then refined NurtureBot v1’s prompt architecture and performed red-team testing (Hong et al., 2024), by involving the two project managers of the original Nurture project, who had previously designed, facilitated, and trained the volunteer moms. Our research team then independently crafted and iteratively refined the prompt.
NurtureBot v1 was designed to replicate the supportive environment of the original peer-to-peer texting interactions while introducing the scalability and accessibility of AI. The chatbot was refined to deliver three key features: empathetic chatting, wellbeing exercises, and parenting information, all adapted for AI-driven interactions. Additionally, it incorporated “conversation starters” thoughtfully crafted by Nurture mentors to begin with and was prompted to perform “contextual reasoning” of the user’s situation before making suggestions. Empathetic chatting was intended to provide ‘active listening’ that reflects warmth and understanding. Wellbeing exercises were included to offer practical tools for managing stress and mental wellbeing, similar to those shared by mentors. Lastly, parenting information was integrated to ensure that users could access reliable and evidence-based resources tailored to their needs, much like the information shared in the original peer support sessions. It is important to note that the development of NurtureBot v1 was limited to zero-shot prompting of the pre-trained GPT-4 model (see Figure 13) and a database of curated resources from the Nurture project to guide its responses. No additional training or fine-tuning of the model was conducted, as we intentionally chose not to train the chatbot on human-to-human conversations, which were deeply rooted in the development of interpersonal relationships and emotional bonding between mentors and new mothers.
3.3. NurtureBot v1 Testing and Analysis
@@
3.3.1. Participants and Method
Once we had secured ethical approval via our institutional ethics board [King’s College London Ethics review - MRA-23/24-44973], we recruited 32 parents of children aged 5 or 6 through the Prolific platform (Prolific, 2014). The average age of participants was 41.17 years, with 20 females and 12 males. Five participants (15%) were in the habit of using ChatGPT everyday. Participants self-reported no acute stress or mental or physical health issues during the study period, nor to have received therapy in the 12 months prior. The study’s purpose and methodology were explained to parents through an information sheet, after which informed consent was obtained. They were paid  \$3 for each day of participation.
The decision to include parents of children aged 5 or 6 as participants in our study was a carefully considered methodological choice aimed at balancing ethical considerations and the practicalities of our exploratory work. We consciously selected this ‘slightly out-of-target’ group as they are likely to still vividly recall and relate to challenges from early parenthood, yet are past the most immediate and demanding stages of infant care. This choice minimised potential emotional risks, as participants in the acute perinatal period might find discussions about parenting challenges triggering without adequate real-time support safeguards. Moreover, our approach ensured that the study was not overly time-intensive for participants already navigating complex, daily baby-related routines. In addition, leveraging adjacent populations is an established approach in exploratory and design studies, similar to engaging individuals with prior therapy experience to inform therapeutic AI systems (Veldmeijer et al., 2023). Our research aimed to simulate and reflect upon perinatal experiences rather than measure clinical outcomes such as Generalised Anxiety Disorder (GAD), aligning it with ethical AI principles that discourage unnecessary exposure of vulnerable populations to experimental settings. By recruiting participants with lived perinatal experience, we gained valuable insights while safeguarding the wellbeing of a demographic closely related to our target population. This approach aligns with exploratory HCI methodologies such as proxy design (Islind et al., 2023) and ethical guidelines outlined in digital parenting support studies ((2020), BERA).
Over the five-day period, participants engaged with NurtureBot v1 for at least five minutes each day, with each day dedicated to simulating a specific stage of early parenthood: pregnancy, the week of childbirth, few weeks postpartum, one month, and six months after birth. The objective was to assess how well the NurtureBot v1 could support them in these hypothetical scenarios. Studies on well-being chatbots support the feasibility of short, focused interactions. For example, the 21-Day Stress Detox chatbot demonstrated effectiveness with sessions lasting approximately 5–7 minutes, providing stress management techniques and maintaining user engagement (Schueller et al., 2021). Similarly, a recent study on GPT-3-based mental well-being chatbots designed five-minute interactions to help users manage their mood, further validating the efficacy of short interaction durations in achieving meaningful outcomes (Doe et al., 2022). These examples align with our approach of employing five-minute everyday interactions for five days to capture user insights effectively.
Each day, participants interacted with the chatbot, receiving responses that included links to resources, empathetic messages, and suggestions for wellbeing exercises (see Figure 3). At the end of each day, after interacting with NurtureBot hosted on a website via Streamlit (Streamlit, 2019), participants were asked to complete a survey via Qualtrics (Qualtrics, 2005), providing feedback on their experience with the chatbot by answering the questions listed in Figure 2. On the final day, they also completed the Chatbot Usability Questionnaire (CUQ), a specialised questionnaire designed to access the usability of healthcare chatbots (Holmes et al., 2019).
@@ 3.3.1. Participants and Method
3.3.2. Data Collection and Analysis
Over the course of the five-day study, 144 conversations and associated feedback were generated, denoting a daily-average of 90% (28.8 out of 32 participants). All data was securely stored on our university’s servers, compliant with GDPR (Union, 2016), and no personal data was collected, consistent with Prolific platform’s design.
The qualitative analysis of the feedback from the daily surveys (See questions 1 and 2 in Figure 2) was conducted using a thematic analysis approach (Braun and Clarke, 2012). Our analysis focused on the instances that captured parent reactions to interacting with NurtureBot. The first author conducted the initial analysis using inductive, open coding, and regularly met with the second author to discuss emerging themes and explore alternative interpretations of participant responses. While participants did share positive feedback on several aspects of the chatbot, highlighting its usefulness and relevance, especially during early stages of parenthood, we do not explore the parents perspectives on the potential of LLM-based parental wellbeing support as this is reported in the literature (Ashraf, 2024), and to keep our focus on answering RQ1. Hence, we focused on analysing 131 of the 288 (45.48%) pieces of feedback that highlighted negative interactions and suggestions for improvement. This led us to the identification of 28 unique pain points, which were further grouped into seven overarching themes, converged in three subsequent meetings.